5.20-90)

Please show steps. Just confused on how to calculate. Thanks!
Answer #1

- for 5. > گ ( 1 ( a a ) ) 1 - . ( یارم ! ہے

So, in your case, λ = 90, μ = 90 and s = 5.

IME

= (1/0!) * 4.5^0 + (1/1!) * 4.5^1 + (1/2!) * 4.5^2 + (1/3!) * 4.5^3 + (1/4!) * 4.5^4

= (1/1) * 4.5^0 + (1/1) * 4.5^1 + (1/2) * 4.5^2 + (1/6) * 4.5^3 + (1/24) * 4.5^4

= 47.8984

د (2)1 -ي (رای

= (1/5!) * (4.5)^5 * 5*20/(5*20 - 90)

= (1/120) * (4.5)^5 * 5*20/(5*20 - 90)

= 153.773

So,

P0 = 1 / (47.8984 + 153.773) = 0.004959